What is the length of the arc on a circle with radius 16 ft intercepted by a 60° angle?


a) 5.3


b) 15.3


c) 16.8


d) 960
1
Expert's answer
2018-11-26T16:39:13-0500

Answer on Question #83354, Math/ Trigonometry

What is the length of the arc on a circle with radius 16 ft intercepted by a 60° angle?


a) 5.3


b) 15.3


c) 16.8


d) 960

Solution

S= r*Θ

60 = 60* (π/180)= π/3

S=16*(π/3) = 16π/3 = 16.8

Answer: c) 16.8